OpenStudy (anonymous):

Prove that (1 + x)^n > or equal to (1 + nx) , for all natural number n , where x>-1??

mathslover (mathslover):

Restrict x > -1, so 1+x > 0. - For n = 1, obviously (1+x)1 = (1+x) = 1 + nx. - Suppose true for n. Then (1+x)n+1 = (1+x)n(1+x) >= (1+nx)(1+x) since we're supposing the statement true for n and since (1+x) > 0 = 1 + nx + x + nx2 by multiplication >= 1 + nx + x since x2 >= 0 = 1 + (n+1)x QED. Note that substituting for a multiplicative quantity in inequalities only works when all elements are positive, so we need (1+x) > 0

OpenStudy (anonymous):

hey it is Mathamatecial Induction Problem

OpenStudy (anonymous):

@ujjwal @UnkleRhaukus @Callisto @myininaya @heena @eseidl @quarkine

OpenStudy (anonymous):

(1+x)^n>1+nx

OpenStudy (anonymous):

((1+x)^n)(1+x)>(1+nx)(1+x) (1+x)^n+1>(1+nx+x+x^2n)>1+(n+1)x.........

OpenStudy (anonymous):

how????

OpenStudy (anonymous):

see (1+x)^n>1+nx............u assume this r8

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

then multiply both side with (1+x)

OpenStudy (anonymous):

since x>-1 sign doesnt change

OpenStudy (anonymous):

OK

OpenStudy (phi):

mathlover's answer is induction. looks ok except for the lousy syntax, which makes it hard to decipher.

OpenStudy (anonymous):

yea...mathlover already gave the answer.!!

OpenStudy (anonymous):

but i did nt understand plzzz @A.Avinash_Goutham plzz continoue

OpenStudy (anonymous):

soo LHS is (1+x)^(n+1)............

OpenStudy (anonymous):